Could somebody help me with this problem?
sqrt(3x+4)-1 = sqrt(2x+1)
Thanks

Answer by venugopalramana(3286)


SEE THE FOLLOWING EXAMPLE AND DO BY SAME METHOD.IF YOU STILL HAVE DIFFICULTY COME BACK
sqrt(13-x) = x-1
square both sides
13-x=(x-1)^2=x^2+1-2x
x^2+1-2x-13+x=0
x^2-x-12=0
find the product of coefficients of x^2 term(1 here)and constant term(-12 here)=-12..now find 2 factors for this whose sum is coefficient of x term(-1 here)..they are -4 and +3...so
x^2-4x+3x-12=0
x(x-4)+3(x-4)=0
(x-4)(x+3)=0
so x=4 or -3